OpenStudy (anonymous):

I WILL MEDAL!! Fifteen is 5% of what number?

OpenStudy (anonymous):

Is it 100

OpenStudy (alexandervonhumboldt2):

x*5/100=15

OpenStudy (alexandervonhumboldt2):

x=15*20

OpenStudy (mimi_x3):

ummm just so you shld know 5% is equal to 5/100 or .05

OpenStudy (anonymous):

In order to find out hat 15 is 5% of what number, Lets assume that the number is x. Therefore, we form an equation: 15=5% of x 15=(5/100) *x Taking 5/100 on the other side: 15*100/5=x or, x=15*100/5 x= 1500/5 x= 300 Therefore, we find out that x is equal to 300, which means that 15 is 5 percent of 300.